# The Experimental Unit Information Index: Balancing Evidentiary Value and Sample Size of Adaptive Designs

### v1

@@ 4.2 N-hacking and constrained sample augmentation
The values of EUII shown in Figure 4 (first-order) and 5 (second-order) are based on simulations for different values of $n_{\scriptsize\texttt{Max}}$, the prior probability $\Pr(H_{1})$ and the true standardized mean difference (SMD). In each Figure, the top panel shows methods which cannot stop early for futility, whereas the bottom plot shows methods which can also stop for futility (including the method from Reinagel). Additional stopping rules for futility has been reported to lead to the saving of resources of up to 30% compared to standard group-sequential methods (Neumann et al., 2017). Here we allow for an additional stop for futility for the group-sequential methods based on the concept of predictive power (PP) (Spiegelhalter et al., 1986; Rufibach et al., 2016). Specifically, we stop the experiment for futility, if the preditive power to achieve significance at the next interim analysis (with four additional observations) is below 30%. The predictive power approach takes into account the uncertainty of the current effect estimate in contrast to standard conditional power calculations (Micheloud and Held, 2022). A threshold on predictive power translates to thresholds on the $p$-values that depend on the interim sample size, as shown in Figure 6. A threshold of 30% leads to $p$-value thresholds between 0.1 and 0.2, whereas smaller thresholds for predictive power (10 or 20%) give larger $p$-values thresholds. We also add futility stopping based on predictive power to $N$-hacking to investigate the effect of a varying $p$-value threshold for futility compared to Reinagel’s approach, where the futility threshold is always 0.1.
@@
The results with the second-order approximation in Figure 5 are qualitatively very similar to the first-order approxmiation shown in Figure 4, with slightly higher values of EUII throughout. The underlying values of $\operatorname{\mathsf{E}}(N_{\bullet})$ and $\mathsf{CV}(N_{\bullet})$ for significant and nonsignificant results are shown in Figure 7 and 8. It can be seen that the expected sample size is smaller for methods that allow for futility stopping, but this is not necessarily the case for the coefficient of variation. Without futility stopping, the coefficient of variation of significant findings can take values up to 0.5, if $n_{\scriptsize\texttt{Max}}$ are large.
@@
Comparing the group-sequential methods without futility stopping (top panels) to the approach from Reinagel (dark-orange line in the bottom panels), it is interesting that the approach by Reinagel often performs better in terms of EUII. However, one may argue that this comparison is unfair, because only Reinagel’s “constrained sample augmentation” approach has an additional stop for futility. The bottom panels of Figure 4 and 5 display the resulting EUIIs with futility stopping. The effect of the prior probability on EUII is now somewhat larger, in particular for Haybittle-Peto and O’Brien-Fleming. The EUIIs of the group-sequential methods with additional PP futility stopping increase, with the Pocock or the Haybittle-Peto method now performing best. Pocock has fairly constant values of EUII and is always among the top two methods. Most other methods also have a fairly constant EUII with varying $n_{\scriptsize\texttt{Max}}$, with the exception of Haybittle-Peto, which is increasing from $n_{\scriptsize\texttt{Max}}=16$ to $n_{\scriptsize\texttt{Max}}=32$. As a result, Haybittle-Peto is even better than Pocock if the assumed SMD is small or $n_{\scriptsize\texttt{Max}}$ is large, but can be worse than all other methods if $n_{\scriptsize\texttt{Max}}$ is small. O’Brien-Fleming is better than Reinagel for a small effect size, but for a large effect size it is the other way round. Finally $N$-hacking plus predictive power is very similar to Reinagel’s approach, slightly better for $\mbox{SMD}=1$ and slightly worse otherwise.
@@
Figure 7: Expected sample sizes with significant (top) and nonsignificant (bottom) results of several group-sequential methods without futility stopping and with futility stopping.
Figure 8: Coefficient of variation of sample sizes with significant (top) and nonsignificant (bottom) results of several group-sequential methods without futility stopping and with futility stopping.
@@ 5 Discussion
We have introduced a new measure to quantify the evidentiary value of one experimental unit, the experimental unit information index. Properties in standard statistical test procedures have been described and the corresponding asymptotic values have been derived. Two application have shown the benefits of this measure to compare more complex statistical procedures with efficacy and futility stopping. It would be interesting to investigate more systematically the optimal choice for efficacy and futility stopping in terms of EUII. In the adaptive designs considered in Sections 4.1 and 4.2, we have used a standard overall significance level of 2.5% (one-sided) and 5% (two-sided) to determine the nominal efficacy thresholds of the different adaptive designs considered. Other overall significance values may lead to different (and perhaps larger) values of EUII.
@@
Code and data to reproduce our results are openly available at https://doi.org/10.5281/zenodo.17671982. We used the statistical programming language R version 4.4.2 (2024-10-31) for analyses (R Core Team, 2024) along with the xtable (Dahl et al., 2019), biostatUZH (Held et al., 2024), ReplicationSuccess (Held, 2020), ggplot2 (Wickham, 2016), cowplot (Wilke, 2024), ggpubr (Kassambara, 2023), dplyr (Wickham et al., 2022), tidyr (Wickham et al., 2024), rpact (Wassmer and Brannath, 2016), gsdesign (Anderson, 2024), SimDesign (Chalmers and Adkins, 2020), and knitr (Xie, 2024) packages.
@@ Acknowledgments
We appreciate helpful comments by Tim Friede, Florian Frommlet and Martin Posch.
@@ Appendix A Asymptotic considerations
here $\Phi(.)$ denotes the standard normal cdf, $\delta>0$ is the assumed standardized effect size of interest and $\alpha$ is the significance level from above. For a two-sample $z$-test with total sample size $n$, the power is
@@ B.2 Data-generating mechanism
The data-generating mechanism (DGM) closely resembled the DGM from the simulation study by Reinagel (2023). In each simulation repetition, a data set consisting of $n_{\text{Max}}$ observations per group is simulated. The observations in the control group are simulation from a standard normal distribution whereas the observations in the treatment group are simulated from a normal distribution with mean $\delta$ and variance 1. The parameter $\delta$ thus represents the standardized mean difference between treatment and control. Similar to Reinagel, the maximum sample size $n_{\text{Max}}$ is varied from 12 to 32 in steps of 4, representig realistic sample sizes for animal experiments. The standardized mean difference $\delta$ is varied from 0 to 1 to represent no (0), small (0.3), medium (0.5), large (0.8), and very large (1) effect sizes. Both factors are varied in a fully factorial way, leading to $6\times 5=30$ simulation conditions, see Table 4.
@@
Standardized mean difference $\delta\in\{0,0.3,0.5,0.8,1\}$
@@ B.3 Estimands and other targets
The estimand of interest is the standardized mean difference $\delta$. The target of interest is the null hypothesis $H_{0}\colon\delta=0$.
@@ B.4 Methods
$\mathrm{PP}_{i+1}=\Phi\left(-t_{1-a_{i+1}/2,2n_{i+1}-2}\sqrt{\frac{1}{1-f}}+z_{i}\sqrt{\frac{f}{1-f}}\right)+\Phi\left(-t_{1-a_{i+1}/2,2n_{i+1}-2}\sqrt{\frac{1}{1-f}}-z_{i}\sqrt{\frac{f}{1-f}}\right)$

### v2

@@ 4.2 N-hacking and constrained sample augmentation
The values of EUII shown in Figure 5 (first-order) and 6 (second-order) are based on simulations for different values of $n_{\scriptsize\texttt{Max}}$, the prior probability $\Pr(H_{1})$ and the true standardised mean difference (SMD). In each Figure, the top panel shows methods which cannot stop early for futility, whereas the bottom plot shows methods which can also stop for futility (including the method from Reinagel). Additional stopping rules for futility has been reported to lead to the saving of resources of up to 30% compared to standard group-sequential methods (Neumann et al., 2017). Here we allow for an additional stop for futility for the group-sequential methods based on the concept of predictive power (PP) (Spiegelhalter et al., 1986; Rufibach et al., 2016). Specifically, we stop the experiment for futility, if the predictive power to achieve significance at the next interim analysis (with four additional observations) is below 30%. The predictive power approach takes into account the uncertainty of the current effect estimate in contrast to standard conditional power calculations (Micheloud and Held, 2022). A threshold on predictive power translates to thresholds on the $p$-values that depend on the interim sample size, as shown in Figure 7. A threshold of 30% leads to $p$-value thresholds between 0.1 and 0.2, whereas smaller thresholds for predictive power (10 or 20%) give larger $p$-values thresholds. We also add futility stopping based on predictive power to $N$-hacking to investigate the effect of a varying $p$-value threshold for futility compared to Reinagel’s approach, where the futility threshold is always 0.1.
@@
The results with the second-order approximation in Figure 6 are qualitatively very similar to the first-order approximation shown in Figure 5, with slightly higher values of EUII throughout. The underlying values of $\operatorname{\mathsf{E}}(N_{\bullet})$ and $\mathsf{CV}(N_{\bullet})$ for significant and nonsignificant results are shown in Figures B1 and B2 in Appendix B. It can be seen that the expected sample size is smaller for methods that allow for futility stopping, but this is not necessarily the case for the coefficient of variation. Without futility stopping, the coefficient of variation of significant findings can take values up to 0.5, if $n_{\scriptsize\texttt{Max}}$ are large.
@@
Comparing the group-sequential methods without futility stopping (top panels) to the approach from Reinagel (dark-orange line in the bottom panels), it is interesting that the approach by Reinagel often performs better in terms of EUII. However, one may argue that this comparison is unfair, because only Reinagel’s “constrained sample augmentation” approach has an additional stop for futility. The bottom panels of Figure 5 and 6 display the resulting EUIIs with futility stopping. The effect of the prior probability on EUII is now somewhat larger, in particular for Haybittle-Peto and O’Brien-Fleming. The EUIIs of the group-sequential methods with additional PP futility stopping increase, with the Pocock or the Haybittle-Peto method now performing best. Pocock has fairly constant values of EUII and is always among the top two methods. Most other methods also have a fairly constant EUII with varying $n_{\scriptsize\texttt{Max}}$, with the exception of Haybittle-Peto, which is increasing from $n_{\scriptsize\texttt{Max}}=16$ to $n_{\scriptsize\texttt{Max}}=32$. As a result, Haybittle-Peto is even better than Pocock if the assumed SMD is small or $n_{\scriptsize\texttt{Max}}$ is large, but can be worse than all other methods if $n_{\scriptsize\texttt{Max}}$ is small. O’Brien-Fleming is better than Reinagel for a small effect size, but for a large effect size it is the other way round. Finally $N$-hacking plus predictive power is very similar to Reinagel’s approach, slightly better for $\mbox{SMD}=1$ and slightly worse otherwise.
@@
5 Application: How many animals could be saved with group-sequential designs?
We will now revisit the 2738 animal experiments in neuroscience and metabolism from Bonapersona et al. (2021). We will assess how many animals could have been saved had each experiment also featured an interim analysis. We will also investigate the rejection rates of different group-sequential designs with and without additional stops for futility.
Assume that the interim analysis takes places after the data from $n_{1}$ animals have been collected, whereas the final analysis takes place after $n_{2}>n_{1}$ animals. The corresponding z-values have the marginal distributions $Z_{i}\sim\mathrm{N}(\delta\sqrt{n_{i}},1)$ for $i\in\{1,2\}$, while their correlation is $\mathrm{Cor}(Z_{1},Z_{2})=\sqrt{n_{1}/n_{2}}$ (Jennison and Turnbull, 1999). For one-group comparisons, $n_{1}$ and $n_{2}$ are the actual sample sizes at each stage, but for two-group comparisons, as in the Bonapersona et al. (2021) data, they are the effective sample sizes given by half the harmonic mean of the sample sizes in the control and treatment groups. In the data from Bonapersona et al. (2021), we only know the observed value of $z_{2}$ at the final analysis, but can derive the conditional distribution of $Z_{1}$ given the observed $z_{2}$, which is
$Z_{1}\mid Z_{2}=z_{2}\sim\mathrm{N}\left(z_{2}\sqrt{n_{1}/n_{2}},1-n_{1}/n_{2}\right).$ (23)
Importantly, the standardised effect $\delta$ cancels out, so the conditional distribution of $Z_{1}$ given $Z_{2}=z_{2}$ is the same regardless of whether data are generated under the null $H_{0}\colon\delta=0$ or the alternative $H_{1}\colon\delta\neq 0$. We will now use this result and simulate interim analyses to assess the effect of early stopping.
Table 4: Mean sample size and rejection rate when introducing an additional interim analyses in 2’738 animal experiments in neuroscience and metabolism (Bonapersona et al., 2021). Values represent medians with 2.5% and 97.5% quantiles from 10’000 simulated interim analyses. Methods with futility stopping halt the study if the predictive power for rejecting $H_{0}$ in the final analysis is less than 10%.
Method Mean $n$ $H_{0}$ rejections (%) Interim efficacy stops (%) Interim futility stops (%) Animals saved
No interim analysis 22.5 27.8 0 0 0
Haybittle-Peto 21.2 (21.1, 21.3) 28.3 (28, 28.5) 11.8 (11, 12.5) 0 (0, 0) 3’804 (3’515, 4’099)
O’Brien-Fleming 21.4 (21.3, 21.5) 27.6 (27.4, 27.8) 9.8 (9.1, 10.4) 0 (0, 0) 3’152 (2’891, 3’426)
Pocock 20.6 (20.5, 20.7) 25.9 (25.3, 26.4) 16.7 (15.9, 17.6) 0 (0, 0) 5’384 (5’043, 5’738)
Haybittle-Peto + futility 18.2 (18, 18.4) 27.6 (27.2, 28) 11.8 (11, 12.5) 28.7 (27.2, 30.2) 11’858 (11’263, 12’470)
O’Brien-Fleming + futility 18.4 (18.2, 18.6) 26.9 (26.6, 27.2) 9.8 (9.1, 10.4) 29.4 (27.9, 31) 11’423 (10’832, 12’024)
Pocock + futility 16.7 (16.5, 17) 25.3 (24.7, 25.9) 16.7 (15.9, 17.6) 37.1 (35.6, 38.7) 15’905 (15’286, 16’541)
Table 4 shows results from simulating 10’000 interim $z$-statistics $Z_{1}$ from (23) for each observed final $z$-statistic $z_{2}$ in the Bonapersona et al. (2021) data set. The interim analyses were conducted halfway through the experiment, where the the number of animals in control and treatment groups has been divided by two and rounded to the next integer, if non-integer values were obtained. We considered Haybittle-Peto, O’Brien-Fleming, and Pocock stopping rules, both with and without possibility for futility stopping. Following the recommendation by Matthews (2006), we use a nominal interim significance threshold of 0.01 (rather than 0.001) for the Haybittle-Peto method with just one interim analysis. This results in a further inflated Type-I error rate, whereas the other two group-sequential methods control the Type-I error rate at the chosen two-sided $\alpha=5\%$ level. We have used a futility stopping threshold of 10% for the predictive power at the end of the experiment.
Without futility stopping, the mean sample size is reduced by around 1-2 animals per experiment compared to no interim analysis. However, further reductions can be achieved with additional futility stopping. The Pocock design with futility stopping saves nearly 6 animals per experiment, totalling to around 16’000 animals. This is accompanied by a drop in $H_{0}$ rejection rates from 27.8% to 25.3%. It should be noted that reduced rejection rates of adaptive designs are not necessarily a disadvantage. In the Bonapersona et al. (2021) data set, the ground truth is unknown, and as such we cannot know whether a $H_{0}$ rejection is a true or false positive. Note that the Haybittle-Peto design has even higher rejection rates than without an interim analysis because of its inflated Type-I error rate.
Interestingly, 37.1% of experiments stop for futility with the Pocock design, whereas this number is below 30% for the other two group-sequential designs due to their larger nominal significance thresholds at the final analysis. Stops for efficacy are also more common with the Pocock design, where 1 in 6 experiments could have been stopped early for efficacy, compared to 1 in 8 with the Haybittle-Peto design and 1 in 10 with O’Brien-Fleming. In summary, using group-sequential designs that allow for futility stops could save a large number of resources and animals. In particular, the Pocock method combined with futility stopping results in the largest savings in sample size, with only modest reductions in $H_{0}$ rejections. This is consistent with the findings from our simulation study described in Section 4.2, where Pocock emerged as the best method in terms of EUII, admittedly in a setting with several (rather than one) interim analyses.
@@ 6 Discussion
We have introduced a new measure to quantify the evidentiary value of one experimental unit, the experimental unit information index. Properties in standard statistical test procedures have been described and the corresponding asymptotic values have been derived. Two applications have shown the benefits of this measure to compare more complex statistical procedures with efficacy and futility stopping. It would be interesting to investigate more systematically the optimal choice for efficacy and futility stopping in terms of EUII. In the adaptive designs considered in Sections 4.1 and 4.2, we have used a standard overall significance level of 2.5% (one-sided) and 5% (two-sided) to determine the nominal efficacy thresholds of the different adaptive designs considered. Other overall significance values may lead to different (and perhaps larger) values of EUII.
@@
The CC-By 4.0 licensed data set from Bonapersona et al. (2021) was downloaded from https://doi.org/10.17605/OSF.IO/WVS7M. Code and data to reproduce our results are openly available at https://doi.org/10.5281/zenodo.17671982. We used the statistical programming language R version 4.4.2 (2024-10-31) for analyses (R Core Team, 2024) along with the xtable (Dahl et al., 2019), biostatUZH (Held et al., 2024), ReplicationSuccess (Held, 2020), ggplot2 (Wickham, 2016), cowplot (Wilke, 2024), ggpubr (Kassambara, 2023), dplyr (Wickham et al., 2022), tidyr (Wickham et al., 2024), rpact (Wassmer and Brannath, 2016), gsdesign (Anderson, 2024), SimDesign (Chalmers and Adkins, 2020), and knitr (Xie, 2024) packages.
@@ Acknowledgments
We appreciate helpful comments by Tim Friede, Florian Frommlet and Martin Posch. We thank Bonapersona et al. (2021) for openly sharing their data.
@@ Appendix A Asymptotic considerations
here $\Phi(.)$ denotes the standard normal cdf, $\delta>0$ is the assumed standardised effect size of interest and $\alpha$ is the significance level from above. For a two-sample $z$-test with total sample size $n$, the power is
@@ B.2 Data-generating mechanism
The data-generating mechanism (DGM) closely resembled the DGM from the simulation study by Reinagel (2023). In each simulation repetition, a data set consisting of $n_{\text{Max}}$ observations per group is simulated. The observations in the control group are simulation from a standard normal distribution whereas the observations in the treatment group are simulated from a normal distribution with mean $\delta$ and variance 1. The parameter $\delta$ thus represents the standardised mean difference between treatment and control. Similar to Reinagel, the maximum sample size $n_{\text{Max}}$ is varied from 12 to 32 in steps of 4, representig realistic sample sizes for animal experiments. The standardised mean difference $\delta$ is varied from 0 to 1 to represent no (0), small (0.3), medium (0.5), large (0.8), and very large (1) effect sizes. Both factors are varied in a fully factorial way, leading to $6\times 5=30$ simulation conditions, see Table 5.
@@
Standardised mean difference $\delta\in\{0,0.3,0.5,0.8,1\}$
@@ B.3 Estimands and other targets
The estimand of interest is the standardised mean difference $\delta$. The target of interest is the null hypothesis $H_{0}\colon\delta=0$.
@@ B.4 Methods
$\mathrm{PP}_{i+1}=\Phi\left(-\frac{t_{1-a_{i+1}/2,2n_{i+1}-2}}{\sqrt{1/f-1}}+\frac{z_{i}}{\sqrt{1-f}}\right)+\Phi\left(-\frac{t_{1-a_{i+1}/2,2n_{i+1}-2}}{\sqrt{1/f-1}}-\frac{z_{i}}{\sqrt{1-f}}\right)$
@@ B.5 Performance measures
B.6 Additional results
Figures B1 and B2 shows the means and coefficient of variations of sample sizes with significant and nonsignificant results.
Figure B1: Expected sample sizes with significant (top) and nonsignificant (bottom) results of several group-sequential methods without futility stopping and with futility stopping.
Figure B2: Coefficient of variation of sample sizes with significant (top) and nonsignificant (bottom) results of several group-sequential methods without futility stopping and with futility stopping.